OpenStudy (anonymous):

Lucy's penny bank is 2/3 full . After she takes out 80 pennies it is 1/2 full . How many pennies can her bank hold? @kelliegirl33

OpenStudy (anonymous):

2/3 - 1/2 = 4/6 - 3/6 = 1/6 80 pennies is 1/6 80/(1/6) = x/(6/6) 1/6x = 80 x = 80/(1/6) x = 80 * 6/1 x = 480 pennies